The human eye contains a molecule called 11

-

cis

-

retinal that changes shape when

struck with

light of sufficient energy, which results in an electrical signal being sent

to the brain. The minimum energy required to change the conformation of 11

-

cis

-

retinal within the eye is 164 kJ/mol. Calculate the

longest wavelength (in nm)

visible to the human

eye.
Answer #1


Given:
Energy of 1 mol = 1.64*10^2 KJ/mol
= 1.64*10^5 J/mol
Find energy of 1 photon first
Energy of 1 photon = energy of 1 mol/Avogadro's number
= 1.64*10^5/(6.022*10^23)
= 2.723*10^-19 J
This is energy of 1 photon

use:
E = h*c/lambda
2.723*10^-19J =(6.626*10^-34 J.s)*(3.0*10^8 m/s)/lambda
lambda = 7.30*10^-7 m

= 730 nm
Answer: 730 nm